Make New Entry,
Make Followup Entry
User name Bob / Ramesh
Log entry time 13:00:25 on November23,2009
Entry number 302075
keyword=L-HRS evtype bits are fine
The L-HRS evtype bits are fine, and have been since the start
of the experiment. Just they were not decoded correctly.
Ramesh and I fixed this in THaDecData.C and in db_D.dat.
The data are in channels 0 - 11 of roc 4 slot 13.
We also see the unprescaled data in the higher channels
(I think it's in chan 12-23 but needs to be checked).
The cut on the TDC data is 200 to 1500 which is pretty
wide (can cause "background"), and the peak seems to be
around 1000. This can be adjusted later.